When two numbers x and y are added together, the sum is 52. y is 3 times larger than x. Find y.
Answer by Alan3354(69443)

When two numbers x and y are added together, the sum is 52. y is 3 times larger than x. Find y.
------------
y = x + 3 times x = 4x
x + y = 52
x + 4x = 52
x = 10.4
y = 4x
y = 41.6
==============
3 times larger = 4 times as large, not 3 times as large.